OpenStudy (anonymous):

Find the inverse of the function \[f(x)=\sqrt{x}+6\]

OpenStudy (anonymous):

Rearrange so x=f(x)/6 then invert f^-1(x)=x/6

OpenStudy (bahrom7893):

Wrong

OpenStudy (bahrom7893):

replace f(x) by x and x by f(x) and solve for f(x)

OpenStudy (bahrom7893):

x=sqrt(f(x)) + 6

OpenStudy (bahrom7893):

Solve for f(x)

OpenStudy (anonymous):

An inverse function in this case will be something that turns \[\sqrt{x}+6 \rightarrow x\]

OpenStudy (bahrom7893):

x-6=sqrt(f(x)) inverse is: f^-1(x) = (x-6)^2

OpenStudy (anonymous):

In comparison to the normal function:\[x \rightarrow \sqrt{x}+6\]

OpenStudy (anonymous):

@bahrom7893 does this method work for finding all inverse functions?

OpenStudy (bahrom7893):

well i've always used it. Idk of any exceptions.

OpenStudy (bahrom7893):

http://www.purplemath.com/modules/invrsfcn3.htm

OpenStudy (bahrom7893):

oh it does... but i showed u the steps in reverse...

OpenStudy (bahrom7893):

When you have an equation, solve for x. You'll get: x = y+sqrt(y) or something.. <-- I made that up, so the inverse will be: f^-1(x) = x+sqrt(x)

OpenStudy (anonymous):

your hella cool. thanks.

OpenStudy (anonymous):

let y= sqrt(x)+6 4y-6=sqrt(x) x=(4y-6)^2 replace x by inverse (x) and y by x, therefore inverse(x)= (4x-6)^2